# A Quantitative Evaluation of the Expressivity of BMI, Pose and Gender in Body Embeddings for Recognition and Identification

### v1

@@ III-C Expressivity of Body Features
IV EXPERIMENTS
IV-A Dataset and Settings
Figure 4: Attribute annotated exemplar images from the BRIAR dataset.
We use the BRIAR 1-5 dataset [11] which has 382,229 images and 170,522 videos of which we extract 704,999 frames of 2,077 unique identities for our work. In our curated data subset, there are 887 males and 1,190 unique female subjects. BRIAR is a massive unconstrained person re-identification dataset containing over 40,000 videos and 1 million images captured across challenging real-world conditions, including varying clothes, distances (100m, 200m, 400m, 500m, 600m, 800m, 1km), altitudes (close range to UAV), and environmental factors like occlusion, blur, and physical turbulence. The dataset consists of 5 variants (BRIAR-1, 2, 3, 4, and 5), each with increasing complexity in terms of number of identities, distractors, and capturing conditions, making it a comprehensive benchmark for testing re-identification systems in unconstrained scenarios. Some example images from the dataset along with the required attribute annotation is shown in Figure 4.
IV-B Integration of MINE with SemReID
SemReID is a self-supervised person re-identification model that introduces a novel Local Semantic Extraction (LSE) module. This module leverages keypoint predictions to guide an interactive segmentation model, Segment Anything Model (SAM), in extracting precise local semantic masks for various body parts. These masks enable the model to isolate and extract fine-grained biometric features that enhance its ability to discern unique identities. Trained using a teacher-student framework, the SemReID model incorporates multiple loss functions to ensure the learned representations are robust and transferable. This approach simplifies the inference process by utilizing only the teacher network, paired with a single linear layer, to perform efficient re-identification. During inference, the model processes input data—including both still images and video frames—through the teacher encoder. By concatenating global and local features, the SemReID model effectively generalizes across diverse re-identification domains, negating the need for domain-specific adaptations or fine-tuning.
To integrate MINE into the SemReID pipeline, an auxiliary neural network is employed to estimate MI by maximizing the Donsker-Varadhan (DV) lower bound. As described earlier, MINE provides a robust framework for estimating MI between learned feature representations and sensitive attributes, such as gender, pose, or identity. The implementation of this integration is detailed in Algorithm 1, which outlines the process for computing the expressivity of the learned representations.
@@ IV-B Integration of MINE with SemReID
The process begins with a dataset consisting of $n$ images, each associated with a corresponding sensitive attribute vector $\mathbf{A}\in\mathbb{R}^{n\times 1}$. Features for these images are extracted from a specific layer $L$ of the SemReID model after a training epoch. These features, denoted as $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$, represent both global and local information captured by the model. To prepare the input for the MINE network, the features and attributes are concatenated to form an augmented matrix $\mathbf{X}=[\mathbf{F}|\mathbf{A}]$. The MINE network $T_{\theta}$ is then initialized based on the dimensions of $X$. The network architecture, described in Figure 2, is a multi-layer perceptron (MLP) comprising two hidden layers with 256 and 128 units, respectively. These layers use Exponential Linear Unit (ELU) activations, chosen for their ability to ensure stable training and mitigate issues such as vanishing gradients. The MLP computes $T_{\theta}$, which is used to estimate the MI by maximizing the DV lower bound. To compute the expressivity of the learned representations, the process iterates for $M$ iterations. In each iteration, the MINE network processes the augmented matrix $\mathbf{X}$ to compute an expressivity score $e$. This score reflects the MI between the feature representations and the sensitive attributes. The scores from all iterations are stored in a list $E$, and the final expressivity measure is obtained by averaging these scores, as described in Algorithm 1.
This integration of MINE with SemReID provides a principled approach to evaluating the expressivity of learned representations. By quantifying the MI between features and sensitive attributes, it becomes possible to assess the extent to which the model captures attribute-relevant information. Furthermore, the iterative nature of the process ensures robust estimates of expressivity while minimizing biases introduced by stochastic factors during training. Importantly, this integration is computationally efficient, as the MINE network operates on features extracted from the SemReID model, making it scalable.
IV-C Hierarchical and Temporal Analysis of Attribute Influence
@@
To achieve a comprehensive end-to-end analysis of attribute influence within our framework, we evaluate feature and attribute correlations both hierarchically across the model’s layers and temporally throughout the training process. The SemReID model employs a Vision Transformer (ViT) backbone with 12 attention layers, offering a powerful architecture for capturing both global and local semantic information.
For the hierarchical analysis, we extract features from multiple attention layers, specifically layers 2, 4, 6, 9, and 12. These layers are chosen to provide a granular understanding of how attribute correlations evolve at different depths of the network. By analyzing the intermediate layers, we gain insights into the progression of feature learning, starting from the lower layers that capture basic patterns and spatial information to the deeper layers where high-level semantic representations emerge. This hierarchical evaluation allows us to observe the model’s ability to encode sensitive attributes at various stages of the feature extraction pipeline and how these representations are refined.
For the temporal analysis, we assess the evolution of feature-attribute correlations across training epochs. Building on prior work that established 11 epochs as the optimal training duration for identity recognition, we extend this analysis to include intermediate epochs. Specifically, we extract features from layers 2, 4, 6, 9, and 12 at training epochs 1, 3, 5, 8, and 11. This temporal sampling strategy allows us to investigate the dynamic changes in feature-attribute correlations as the model progresses through training. By observing earlier epochs, such as epochs 1 and 3, we capture the initial stages of representation learning, where the model begins to encode attribute information. In contrast, later epochs, such as 8 and 11, reveal how the model converges and solidifies its understanding of attributes in its learned representations.
This dual analysis framework examining both the hierarchical layers and temporal progression offers a detailed perspective on how attribute information is processed and encoded within the SemReID model. Hierarchical analysis highlights the architectural contribution of different layers to feature representation, while temporal analysis sheds light on the dynamic learning trajectory of attribute representations over time. Together, these analyses enable us to uncover trends and patterns in attribute encoding, providing valuable insights into the model’s learning process in mitigating or amplifying attribute correlations.
IV-D Implementation Details
We initialize the MINE approximation network based on the input dimensions of the augmented matrix. As outlined earlier, the network is trained to estimate the lower bound of the MI between features $F$ and attribute $A$. For this purpose, we employ a simple multi-layer perceptron (MLP) architecture, illustrated in Figure 2, to compute $T_{\theta}$. The MLP consists of two hidden layers with 256 and 128 units, respectively, followed by ELU activation functions. This architecture is used consistently across all experiments. The MINE network is trained using the Adam optimizer with a learning rate of 0.001 and a batch size of 100. When different sets of features are used for $F$, only the input layer’s dimensions are adjusted to match the feature dimensionality. The network is trained until the loss in Equation (7) converges, and the expressivity is calculated as described in Algorithm 1. For all experiments, we set $M=16$. SemReID uses Vision Transformer (ViT) [14] variants as the backbone. Images are processed at $384\times 128$ resolution with a single forward pass. The inference pipeline uses a dual-stream semantic parsing strategy, where global features are processed at $768$ dimension and local semantic features at $3\times 768$ dimension where $3$ is the number of local areas. In our features they are face, upper body, and lower body. The final local features is the average of all local areas. We implement multi-crop feature augmentation [6, 7] using a cascade of $M=2$ global views and $N=3$ local views, followed by $L=12$ cross-attention layers for feature alignment. The identity embedding is computed through a BN layer, optimized for minimal latency while maintaining discriminative capabilities. We extract the features from this trained network which have a dimension of 1536 and pass it through the MINE block after concatenating it with the attribute vector to estimate MI.
V RESULTS AND DISCUSSIONS
We present our observations on the correlation between features and attributes in the SemReID model using MINE. The discussion is structured into three subsections: the first examines feature-attribute correlations within the hierarchical feedforward pass, while the second analyzes the evolution of these correlations throughout the training process and the third explains the advantages of our method.
TABLE I: Summary of values for Gender, BMI, Yaw, and Pitch across different layers.
Layer Gender BMI Yaw Pitch
Attention Layer 2 $6.57\times 10^{-3}$ 0.005 0.13 0.42
Attention Layer 4 0.003 0.149 0.283 0.523
Attention Layer 6 0.004 0.72 0.73 0.73
Attention Layer 9 0.005 0.77 0.73 0.73
Attention Layer 12 0.03 1 0.10 0.41
@@ V RESULTS AND DISCUSSIONS
Figure 5: Expressivity trend of gender, yaw, pitch and BMI in input image over layer-wise learnt features from SemReID.
V-A Layer-Wise Progression in Feedforward Processing
It is crucial to emphasize that the model was trained exclusively on identity labels, without explicit supervision for BMI, gender, or pose attributes. Given that our architecture employs a ViTbase backbone, it consists of 12 transformer encoder blocks, each containing an attention layer. To analyze the evolution of feature representations, we compute the MINE at various depths—specifically after layers 2, 4, 6, 9, and 12—as summarized in Table 1 and depicted in Figure 5. The key observations are as follows:
  • In the initial layers (e.g., Layer 2), MI between the learned representations and pitch/yaw attributes is relatively high, whereas gender and BMI exhibit negligible expressivity. This suggests that lower layers predominantly encode coarse spatial and geometric features, which are crucial for pose estimation.
  • As we progress deeper into the ViT_base network, BMI expressivity increases rapidly, suggesting that this is a fine grained feature and is important for the model. By Layer 6, all attributes exhibit substantial MI with the learned features. However, gender expressivity remains comparatively lower, implying that it is not significantly encoded in the model’s learned representations.
@@ V-A Layer-Wise Progression in Feedforward Processing
Thus, for the semReID model trained purely on identity labels, BMI emerges as the most correlated attribute, followed by pose, with gender being the least correlated. This conclusion aligns with intuitive human perception, as body recognition inherently relies on a person’s shape and pose as primary cues. From a perceptual standpoint, humans often associate body identity with physical attributes such as body shape, proportions, and posture, making BMI and pose naturally dominant in recognition tasks.
We want to clarify at this stage that, the data processing inequality (DPI) [23] states that for three random variables $P$, $Q$, and $R$ forming a Markov chain $P\rightarrow Q\rightarrow R$, the MI satisfies $MI(P,Q)\geq MI(P,R)$. This principle formalizes the idea that no processing of data can increase MI. In the context of neural networks, let $P$ represent a random variable (e.g., BMI, pose, or identity), and $Q$ and $R$ represent features extracted at different layers of the network, where $R$ is a deterministic function of $Q$ (i.e., $R$ corresponds to a deeper layer). Since $R$ is derived from $Q$, $P$, $Q$, and $R$ form a Markov chain, and the DPI implies that the MI between $P$ and the features cannot increase as we go deeper into the network. However, the expressivity results are not strictly monotonically decreasing, which might seem to contradict the DPI. This apparent inconsistency arises because, in our work, expressivity refers to the alignment between feature representations and a given attribute, rather than MI in the strict sense of information theory. As noted in [14], the feature representations in this context are more closely tied to the predictability of an attribute rather than its information-theoretic content. Therefore, expressivity in this work reflects the strength of the relationship between features and attributes, rather than their theoretical MI. This distinction allows us to interpret the expressivity results as a measure of the model’s ability to capture attribute-specific representations, which can vary across layers.
Figure 6: Expressivity trend of gender, yaw, pitch and BMI in input image over epoch-wise learnt features from SemReID.
TABLE II: Summary of values for Gender, BMI, Yaw, and Pitch across different epochs.
Epoch Gender BMI Yaw Pitch
Epoch 1 0.012 0.87 1.3 0.84
Epoch 3 0.402 0.305 0.009 0.004
Epoch 5 0.004 0.76 $6\times 10^{-3}$ $4\times 10^{-3}$
Epoch 8 0.012 0.80 0.05 0.22
Epoch 11 0.03 1 0.10 0.41
V-B Progression of Training
To understand how attribute-feature correlations evolve over training, we analyze the expressivity of gender, yaw, pitch, and BMI in the final attention layer across different training epochs. This is summarized in Table II and visualized in Figure 6. At each epoch, the feature representations are 1536-dimensional, and attribute-wise expressivity is computed using Algorithm 1. We observe that:
@@ V-B Progression of Training
  • A significant reduction in yaw expressivity is observed, while BMI expressivity remains relatively stable. This suggests that yaw-related variations are gradually suppressed as the model prioritizes more identity-relevant features. Notably, both yaw and gender exhibit minimal correlations with the learned representations, implying that the model is learning to be invariant to these attributes in its deeper layers. Conversely, BMI remains a highly correlated feature, followed by pitch, reinforcing its role as a more identity-relevant attribute.
[…]
These findings indicate that the training process makes feature representations increasingly invariant to certain body attributes, such as yaw and gender, while preserving information related to BMI and to some extent pitch.
V-C Advantages of Expressivity for Person ReID
In this subsection we justify the usage of MINE over other existing methods. The key reasons can be summarized as:

### v3

@@ 4.2 Integration of MINE with ReID Models
We analyze attribute correlation trends over three SoTA person ReID models namely Pose-guided Feature Disentangling (PFD)[43], Dc-former[27] and SemReID [20]. PFD introduces a transformer-based framework for occluded person Re-ID, using pose keypoints to guide feature disentanglement between visible and occluded body regions. It leverages a ViT-base backbone to encode patch-level features and a transformer decoder with learnable semantic tokens aligned to pose-guided regions via a Pose-View Matching module. A pose-guided push loss further encourages discrimination of visible parts while suppressing noisy occluded features for robust identity matching. The DC-Former utilizes a ViT-small backbone to improve person ReID by learning multiple diverse and compact embedding subspaces. It uses multiple class tokens within the Vision Transformer framework to represent distinct embedding spaces. A Self-Diverse Constraint (SDC) is applied to encourage these embeddings to be both diverse and compact, improving the model’s ability to distinguish between similar identities. Additionally, a Dynamic Weight Controller (DWC) balances the importance of each embedding space during training, leading to robust and discriminative feature representations. SemReID is a self-supervised person ReID model that introduces a novel Local Semantic Extraction (LSE) module. This module uses keypoint predictions to guide the Segment Anything Model (SAM), producing precise local semantic masks for various body parts. These masks allow for the extraction of fine-grained biometric features, enhancing identity discrimination. SemReID is trained using a teacher-student framework with multiple loss functions to promote robustness and transferability. At inference, only the teacher network and a single linear layer are used, enabling efficient re-identification by processing input data through the teacher encoder. Global and local features are concatenated to facilitate generalization across domains without domain-specific fine-tuning.
@@
To evaluate feature expressivity, MINE is integrated into these pipelines as an auxiliary neural network estimating mutual information (MI) by maximizing the Donsker-Varadhan (DV) lower bound. Given a dataset of $n$ images with corresponding sensitive attributes $\mathbf{A}\in\mathbb{R}^{n\times 1}$, features are extracted from a specific layer $L$, yielding $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$. These features, capturing both global and local cues, are concatenated with attributes to form the input matrix $\mathbf{X}=[\mathbf{F}\,|\,\mathbf{A}]$. The MINE network $T_{\theta}$, a multi-layer perceptron (MLP) with hidden layers of 512 and 128 units and ELU activations, is initialized based on $\mathbf{X}$’s dimensions. Over $M$ iterations, $T_{\theta}$ processes $\mathbf{X}$ to compute expressivity scores $e$, which estimate the MI between features and attributes. These scores are stored in a list $E$, and the final expressivity is computed as their average (Algorithm 1). This integration provides a principled and scalable approach to quantify attribute-relevant information in learned representations. The iterative MINE process ensures stable and unbiased estimates of expressivity while remaining computationally efficient, as it operates on pre-extracted features from these models.
4.3 Hierarchical and Temporal Analysis of Attribute Influence
@@
To comprehensively analyze attribute influence in our framework, we examine feature–attribute correlations both hierarchically across model layers and temporally over training epochs. The SemReID model and the PFD model use a ViT backbone with 12 attention layers, while the Dc-former model uses 8 attention layers to capture rich global and local semantics.
Hierarchical Analysis: We extract features from layers 2, 4, 6, 9, and 12 for the PFD and SemReID models while features from layers 2,3,4,6 and 8 for the Dc-former to study how attribute correlations evolve with network depth. We consider layers 2,4 to be early 6,9 to be mid and 12 to be late for ViT base while 2,3 to be early, 4,6 to be mid and 8 to be late for ViT small. These layers are selected to provide a fine-grained view of the learning process from early layers capturing basic spatial patterns to deeper layers encoding high-level semantics.
Temporal Analysis: We train each of these models for 11 epochs and to assess how these correlations change during training, we analyze the final layer features at epochs 1, 3, 5, 8, and 11. Early epochs (1, 3) highlight the emergence of attribute encoding, while later epochs (8, 11) illustrate how these representations stabilize as the model converges.
Together, this dual analysis, provides a detailed understanding of how attribute information is processed, encoded, and evolved within these models. It uncovers key trends in the model’s capacity to learn, refine, or suppress sensitive attribute correlations throughout training.
Models Rank 1 5 10 TAR @1% @10%
DC-Former 27.98 54.06 62.28 39.84 84.51
PFD 32.92 55.65 75.73 47.97 71.22
SemReID 34.84 56.95 66.33 54.09 89.03
Table 1: Quantitative comparison between DC-Former, PFD, and SemReID for identification task (Best in bold).
4.4 Implementation Details
We initialize the MINE network based on the input dimensions of the augmented matrix, using a two-layer MLP (512 and 128 units, ELU activations) as seen in Fig. 2 to compute $T_{\theta}$. The weights are initialized using Xavier normal initialization. This setup, consistent across experiments, is trained with Adam (learning rate=0.001, batch size=100) until Equation (7) converges. Only the input layer adapts to the feature dimensionality. Expressivity is calculated per Algorithm 1, with $M=5$. For SemReID, we use ViT variants [14] with $384\times 128$ inputs in a single forward pass. A dual-stream setup extracts 768-dim global and $3\times 768$-dim local semantic features (face, upper, lower body), averaged for the final local embedding. Multi-crop augmentation [6, 7] uses $M=2$ global and $N=3$ local views, followed by $L=12$ cross-attention layers. Identity embeddings are computed via a BN layer for efficiency. Final 1536-dim features are concatenated with attribute vectors and fed to MINE to estimate MI. DC-Former [27] uses multiple class tokens with SDC ($\lambda=1$) to create diverse embedding subspaces, using dynamic weight controller for token balancing. PFD [43] utilizes overlapping patch embedding (step size=12), pose-guided feature aggregation via HRNet-estimated keypoints (threshold $\gamma=0.2$), and part view based decoder with $N_{v}=17$ learnable semantic views; decoder layers set to 6.
Gender Yaw BMI Pitch
No. DCFormer SemReID PFD DCFormer SemReID PFD DCFormer SemReID PFD DCFormer SemReID PFD
Layer early (2-s,2-b) 0.3252 0.0978 0.3250 0.3109 0.0026 0.5443 0.3588 0.0950 0.5153 0.4430 0.1461 0.6151
early (3-s, 4-b) 0.3330 0.0999 0.4176 0.3595 0.0210 0.7181 0.4102 0.2174 0.6509 0.4733 0.2243 0.7259
mid (4-s, 6-b) 0.3490 0.1829 0.4321 0.3458 0.0504 0.7611 0.5023 0.2816 0.7940 0.4203 0.2998 0.7678
mid (6-s, 9-b) 0.3101 0.2150 0.4428 0.1909 0.0716 0.5369 0.6101 0.4714 1.2750 0.3440 0.3190 0.5925
late (8-s,12-b) 0.2856 0.1586 0.4280 0.1456 0.0119 0.4074 0.6832 0.5740 1.2316 0.3347 0.2727 0.4950
Epoch 1 0.3272 0.2115 0.4864 0.2454 0.2519 0.7119 0.4968 0.4283 0.9898 0.3065 0.3722 0.7084
3 0.2980 0.1388 0.4403 0.1772 0.000094 0.6426 0.6198 0.3724 1.0660 0.3332 0.2419 0.5998
5 0.2945 0.1224 0.4425 0.1756 0.000029 0.5512 0.6580 0.4194 1.1495 0.3172 0.2052 0.6097
8 0.2914 0.1461 0.4211 0.1896 0.00145 0.4328 0.7047 0.5479 1.2280 0.3342 0.2517 0.5251
11 0.2856 0.1586 0.4280 0.1456 0.0119 0.4074 0.6832 0.5740 1.2316 0.3347 0.2727 0.4950
Table 2: Expressivity scores for Gender, Yaw, BMI, and Pitch across layers and training epochs for DCFormer, SemReID, and PFD. x-s and y-b denote the $x^{th}$ and $y^{th}$ layer of ViT small and base respectively.
5 Results and Discussions
On the person ReID task for BRIAR, SemReID demonstrates the strongest performance (Table 1). In this section, we analyze feature-attribute correlations using MINE. The discussion is structured into three subsections: (1) spatial trends across the hierarchical feedforward pipeline, (2) temporal progression during training, and finally (3) the advantages of our method.
5.1 Hierarchical Feedforward Progression of Attribute Expressivity
@@
Figure 4: Expressivity trends of gender, yaw, pitch and BMI in input image over layer-wise learnt features from SemReID.
To study the evolution of attribute correlations across network depth, we extract features from layers 2, 4, 6, 9, and 12 for the PFD and SemReID models, and from layers 2, 3, 4, 6, and 8 for DCFormer. For ViT-Base (PFD and SemReID), we categorize layers 2 and 4 as early, 6 and 9 as mid-level, and 12 as late. For ViT-Small (DCFormer), layers 2 and 3 are considered early, 4 and 6 as mid-level, and 8 as late. Figure 4 visualizes attribute expressivity per layer, and Table 2 reports corresponding values. Our analysis reveals the following trends:
  • BMI Encodes Deeply Across All Models: BMI expressivity increases consistently with depth in all models. PFD shows the sharpest rise from around 0.6 in early layers to over 1.2 in the final block—highlighting that BMI becomes increasingly entangled in deeper feature hierarchies. This suggests that morphometric cues are emphasized more in later layers.
  • Yaw Peaks in Mid-Layers and Then Drops: Yaw exhibits a peak in the mid-layers and declines in the final block across all models. This is most prominent in PFD, where yaw rises to $\sim 0.75$ mid-network and then drops to $\sim 0.4$. The pattern suggests intermediate layers encode pose cues strongly, which are later partially suppressed as identity features dominate.
  • Pitch Follows a Similar Pattern to Yaw, But Less Drastic: Pitch also peaks at the mid-layer stage before dropping in the late layers. However, unlike yaw, pitch maintains moderate expressivity even in deeper layers, particularly in DCFormer and PFD. This highlights partial retention of pitch as a latent factor.
  • Gender Expressivity is Moderate and Stable: Gender shows moderate expressivity across all layers, particularly in PFD and DCFormer, where its scores remain relatively consistent. While it does not peak as strongly as BMI or pose attributes, it is clearly encoded and not suppressed. In SemReID, gender shows a slight rise in mid-layers before declining in the final block. Overall, gender expressivity is less depth-sensitive but remains a persistent latent factor across the network hierarchy.
  • Model-wise Trends: PFD encodes all attributes more strongly across all layers, indicating greater attribute leakage and less disentanglement. DCFormer exhibits moderate levels of expressivity and clear hierarchical trends. SemReID consistently suppresses attributes more effectively, especially yaw and gender, suggesting more robust identity representation.
@@
5.2 Temporal Progression of Attribute Expressivity
Figure 5: Expressivity trends of gender, yaw, pitch and BMI in input image over epoch-wise learnt features from SemReID.
We analyze the temporal evolution of MINE expressivity scores across epochs for the same four attributes measured at the final attention layer of the three models: SemReID, DCFormer, and PFD. Figure 5 shows the epoch-wise trends, and Table 2 reports the absolute scores. We observe the following key patterns:
@@ 5.2 Temporal Progression of Attribute Expressivity
  • Yaw Suppression Over Time: All models exhibit a notable decline in yaw expressivity with training, dropping from over 0.7 to around 0.4 by epoch 11 in PFD, and collapsing to near zero after epoch 2 in SemReID. This indicates active suppression of yaw as models optimize for pose invariant identity representation.
  • BMI Persistence Across Models: Unlike other attributes, BMI expressivity shows a rising or stable trend for all models. In PFD, it increases from 1.0 to 1.2, while DCFormer and SemReID show comparatively modest gains. This highlights BMI as a persistent entangled attribute, potentially acting as a confounding factor in identity representations.
  • Gender Expressivity is Low and Stable: Gender shows the lowest expressivity throughout training across models except for yaw. In SemReID, it drops early and stabilizes near 0.13, indicating progressive decoupling over time.
  • Pitch Shows Intermediate Trends: Pitch expressivity stabilizes after epoch 3 for all models, with DCFormer consistently encoding more pitch information than SemReID. The stability suggests that pitch is preserved during training but is less dominant than BMI.
  • Model Comparisons: PFD consistently encodes more attribute information than the other two models across all attributes and epochs. In contrast, SemReID exhibits the lowest attribute expressivity overall, indicating more robust identity feature learning with reduced sensitivity to spurious cues like yaw and gender.
These findings reveal that during training, pose-related cues like yaw are actively suppressed, particularly in SemReID. However, body-morphometric attributes such as BMI persist in the feature space, suggesting a need for targeted disentanglement strategies for such persistent confounders.
5.3 Advantages of Expressivity for Person ReID
In this subsection we further justify the usage of MINE over other existing methods. The key reasons are: